OpenStudy (anonymous):

x=t+6 y=t^2-2 Find the equation of the curve along which the particle moves?

OpenStudy (primeralph):

t = x-6, put that in the next equation

OpenStudy (anonymous):

t = x - 6 y = (x - 6)^2 - 2 y = x^2 - 12x + 36 - 2 y = x^2 - 12x + 34

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @pingeye1

OpenStudy (anonymous):

@tcarroll010 thanx a whole lot @primeralph no help at all

OpenStudy (primeralph):

@pingeye1 you're human, you should be able to figure things out.

OpenStudy (anonymous):

@primeralph no need to talk to me i got my question answered

OpenStudy (primeralph):

you should never have mentioned my name in the first place.